### 30 Jan 2024 v1

# Augmenting Replay in World Models for Continual Reinforcement Learning

Abstract
In continual RL, the environment of a reinforcement learning (RL) agent undergoes change. A successful system should appropriately balance the conflicting requirements of retaining agent performance on already learned tasks, stability, whilst learning new tasks, plasticity. The first-in-first-out buffer is commonly used to enhance learning in such settings but requires significant memory. We explore the application of an augmentation to this buffer which alleviates the memory constraints, and use it with a world model model-based reinforcement learning algorithm, to evaluate its effectiveness in facilitating continual learning. We evaluate the effectiveness of our method in Procgen and Atari RL benchmarks and show that the distribution matching augmentation to the replay-buffer used in the context of latent world models can successfully prevent catastrophic forgetting with significantly reduced computational overhead. Yet, we also find such a solution to not be entirely infallible, and other failure modes such as the opposite — lacking plasticity and being unable to learn a new task — to be a potential limitation in continual learning systems.
1 Introduction
Over a lifetime, a person or intelligent system that interacts with a complex environment can expect to acquire new experiences and learn new tasks, all while old experiences and tasks are retained. This requirement to satisfy two orthogonal goals — the acquisition of new knowledge and the retention of old knowledge is the challenge of continual learning (CL) Ring (1994), and the demand for both desiderata presents the stability-plasticity dilemma Carpenter and Grossberg (1987). A system capable of successfully undertaking CL should meet both of these goals. That is, it should be able to replicate its original performance on an earlier task when revisiting it after having learned other tasks, all while not interfering with its ability to gain performance on new tasks.
Neural networks have proved a powerful tool driving the state-of-the-art in games, control and interactive language models Vinyals et al. (2019); OpenAI et al. (2019); Beattie et al. (2016); Andrychowicz et al. (2020); Brown et al. (2020). However, neural networks have been shown to be highly susceptible to forgetting previously learnt knowledge after training on a new task French (1999). This undesirable property of neural networks has proven troublesome in their integration into CL systems, since catastrophe forgetting is the most reported failure mode McCloskey and Cohen (1989); Khetarpal et al. (2022). Catastrophic forgetting occurs when newly acquired skills compromise already learned ones. The opposite of catastrophic forgetting is forward and backward transfer, in which either earlier or later tasks benefit from learning on separate tasks.
Typically, the traditional RL settings focus on solving one individual task that does not change during learning Khetarpal et al. (2022). Where multiple different tasks are required to be learned, one common approach is to learn all simultaneously, rather than sequentially. This avoids catastrophic forgetting as the data distribution of the learning task never changes, but it is not always applicable, as the environments must generate too many samples or because some tasks are unavailable until others have completed. This is the case in continual RL, where new challenges are often experienced one after another. Therefore, a CL system must be able to learn multiple different tasks sequentially. Additionally, tasks may change subtly or silently over time, so a CL system should also be capable without explicit task identifiers Rolnick et al. (2019).
One method that existing CL systems employ in RL is to store experiences of all tasks, so that the system may ‘remember’ and retain knowledge from previous tasks to prevent catastrophic forgetting Lipton et al. (2018). This is done through a replay buffer, often implemented as a generic first-in-first-out (FIFO). Given that the buffer is sufficiently large and that the learning system is capable of using the knowledge stored in the buffer (i.e., the RL algorithm is not a purely on-policy method), such a solution may enable favourable properties of CL such as reduced forgetting and forward and backward transfer. However, the requirement to store all experiences from all tasks, demands very high storage capacity. This is an unfavourable property, as RL is often done at scale and the large volume of samples to store can present a significant engineering challenge, reducing the scalability and accessibility of the algorithm OpenAI et al. (2019).
The inspiration for many replay-based methods comes from Complementary Learning Systems (CLS) Hassabis et al. (2017); Khetarpal et al. (2022), which describes learning in mammalian brains. The hippocampus memorises recent observations and replays them to the neocortex, which is a slow statistical learner. Replay is interleaved with new experiences, thus mitigating catastrophic forgetting. Interestingly, the neocortex is understood to form a ‘world model’ whose purpose is to predict the consequences of our actions Mathis (2023).
The idea of world models has been exploited in RL algorithms (Model Based RL), where a world model is capable of predicting the effects of actions on the environment Ha and Schmidhuber (2018); Hafner et al. (2019, 2020, 2022, 2023). RL world model approaches are an intuitive choice for CL, as they naturally support off-policy learning; retaining the world model’s knowledge of a task is likely sufficient for preserving performance. Although CLS describes replay to a ‘world model’, traditionally replay in RL is used in Model Free RL to improve the policy, rather than a world model directly.
In this paper, we take the most recent development of world model approaches in DreamerV3 Hafner et al. (2023) and explore its ability in CL when used with memory-efficient replay buffers Isele and Cosgun (2018). Hence, we present World Models with Augmented Replay (WMAR) as a promising solution to CL problems faced by world model approaches, while alleviating memory requirements. We evaluate WMAR on a diverse and challenging set of environments from OpenAI Procgen (procedurally generated games) and Atari. These suites pose distinct challenges as to how similar or dissimilar different tasks are. We used Procgen to create similar tasks with ‘shared structure’, whereas Atari games comprise dissimilar tasks ‘without shared structure’. Our analysis is not restricted to catastrophic forgetting but includes forgetting and backward/forward transfer of skills between tasks.
For a background in RL and CL, including desiderata in CL solutions, see Section S.LABEL:S-sec:background. For Related Work, see Section S.LABEL:S-sec:related_work; including methods to counteract catastrophic forgetting in CL Kirkpatrick et al. (2017); Schwarz et al. (2018); Shin et al. (2017) and continual RL Isele and Cosgun (2018); Rolnick et al. (2019); Kessler et al. (2022).
2 World Models with Augmented Replay (WMAR)
WMAR uses world models and augmented replay buffers to learn tasks sequentially. The world model builds on developments in DreamerV3, and is state-of-the-art in single-GPU RL benchmarks in a number of tasks. The off-policy nature of the RL algorithm allows for the use of replay buffers as a mechanism to prevent forgetting. Furthermore, world models naturally support a memory-based continual learning approach as the strength of the final policy is largely contingent on the quality of the world model’s accuracy in simulating the environment. Hence, preventing forgetting within the world model through any means should retain agent performance in past environments. World models are also task-agnostic and labels specifying the task/environment are not required.
The system consists of the World Model for modelling the environment, an Actor-Critic Controller to act on the environment, and a Replay Buffer. The Replay Buffer is used to train the World Model, and the World Model is able to ‘dream’ experiences, which are used to train the Controller. Each is described below, and more details are found in Section S.LABEL:S-sec:world_models.
The source code is available at https://github.com/cerenaut/wmar.
2.1 World Model
The World Model, see Figure 1, learns a Recurrent State-Space Model (RSSM) Hafner et al. (2019) to predict environment dynamics. At timestep $t$, the RSSM contains a deterministic hidden state $h_{t}$ and stochastic state $z_{t}$, either predicted by output of a variational autoencoder $z_{t}\sim q_{\theta}(z_{t}\mid h_{t},x_{t})$ or by the dynamics predictor from $\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid h_{t})$ in open-loop prediction. RSSM models the stochastic representation of the next state given current and previous observations $x_{1:t}$ and actions $a_{1:t}$. It achieves this through a recurrent layer in the GRU, predicting the deterministic state $h_{t+1}=\text{GRU}(h_{t},z_{t},a_{t})$ and the stochastic state $\hat{z}_{t+1}=f(h_{t+1})$.
Following DreamerV3, the stochastic state $z_{t}$ consists of 32 discrete stochastic units with 32 categorical classes each. Should the RSSM be able to accurately predict the environment dynamics (i.e., $\hat{z}_{t}\approx z_{t}$), it follows that the deterministic and stochastic states together form a Markovian representation of the undisclosed environment state. During open-loop prediction, stochastic state posteriors $z_{t}$ do not exist and the priors are taken for the posteriors instead (i.e. $z_{t}=\hat{z}_{t}$). We use the standard implementation of the GRU with Tanh activation, as we did not observe a significant improvement at the cost of extended wall time.
At timestep $t$, the model state of the trained world model is a Markovian representation of the environment and is the concatenation of the deterministic state $h_{t}$ and stochastic state $z_{t}$. The world model is trained by reconstructing input images, rewards and episode continuation flags ($c_{t}\in\{0,1\}$ depending on whether the episode terminates after the observed state) from the model state. The RSSM component is trained through KL balancing Hafner et al. (2022), which allows the dynamics predictor to model the transition between states.
(a) World Model
(b) Flow of data within the RSSM
Figure 1: World Model overview and RSSM details.
2.2 Actor Critic
The actor and critic (Figure S.LABEL:S-fig:actor_critic) are MLPs mapping the model state to actions and value estimates. They are trained entirely on trajectories generated stochastically by the world model, which is referred to as ‘dreaming’ or ‘imagining’. Starting from model states of inputs sampled from the replay buffer, the actor acts within states predicted by the world model, producing a sequence of states, actions, rewards, and continuation flags. We follow DreamerV3 in training the actor and critic on-policy on these imagined trajectories using reinforcement Williams (1992). As such, inevitable changes in distribution within the model state space are not an issue since imagined trajectories are cheap to generate and do not require interaction with the actual environment, allowing this process to be run to convergence of the actor and critic.
2.3 Replay Buffer
A key objective of WMAR is to limit the size of the replay buffers. While Dreamer maintained a single FIFO buffer containing the last $10^{6}$ observations, we empirically choose a size of $2^{18}\approx 262$K observations for each buffer, as it is significantly (about 4×4$\times$4 ×) smaller than that of Dreamer’s, without noticeably costing performance. We also introduce spliced rollouts, which are a simple and sometimes necessary alternative to storing entire episodes, enabling smaller buffer sizes. The buffer types are shown in Figure 2 and described below.
Figure 2: The use of replay within the world model context. LTDM refers to the long-term distribution matching buffer.
2.3.1 Short-Term FIFO Buffer
We maintain a FIFO buffer with a capacity of $2^{18}$ samples containing the most recent rollout observations. The retention of a short-term FIFO buffer follows CLS theory, suggesting that multiple separate structures within a system are required for learning and CL. This buffer also allows the world model to train on all incoming experiences and biases the training samples to those more recently collected, thus improving convergence properties on the current task.
2.3.2 Long-Term Global Distribution Matching (LTDM)
LTDM Buffer
Matching the global training distribution even with limited capacity in the replay buffer has been shown to be a suitable method to reduce catastrophic forgetting Isele and Cosgun (2018). We use a long-term global distribution matching buffer, also with a capacity of $2^{18}$ samples that contain a uniformly random subset of 512 spliced rollouts. This buffer uses the reservoir sampling strategy by assigning a random value for each rollout chunk as a key in a size-limited priority queue, preserving the experiences with the highest key values, and discarding the remainder.
Spliced Rollouts
Restricting the size of the replay buffer can cause it to become full with only a few episodes, especially as episode lengths may be in the thousands. Containing too few unique episodes may cause the training data to become highly unrepresentative of the general environment states, thus reducing the world model’s accuracy in modelling the environment and likely causing a subsequent loss of performance from the actor. This is especially relevant for the long-term global distribution matching buffer, which may store relatively few samples from each environment.
Hence, while typical implementations store complete rollouts, we splice rollouts into smaller chunks of size 512 where necessary, which we find to not adversely affect performance. The implication is that the long-term global distribution matching buffer will operate over the spliced rollouts rather than entire rollouts, thus providing a guarantee on the granularity for its sampling of data.
Remainders of rollouts with fewer than 512 states are concatenated prior to the next episode with an appropriate reset flag indicating the start of a new episode. To improve the efficiency of the rollout process, we truncate episodes after a fixed number of steps, so that the final number of environment steps is identical in every training iteration. We find spliced rollouts to be a simple solution to control granularity in the distribution matching buffer, especially when episodes are extremely long.
2.3.3 Augmented Replay Buffer
We combine the equally sized short-term FIFO buffer $\mathcal{D}_{1}$ with the long-term global distribution matching buffer $\mathcal{D}_{2}$ by using both buffers in parallel. Data from both buffers are uniformly sampled for each training minibatch (Algorithm S.LABEL:S-alg:combined_buffers).
2.4 Task Agnostic Exploration
The exploration of new environments may pose a challenge, as policies that have already been trained on previous tasks may lack the randomness required to adequately explore the state space of a new task. To address this challenge, we use developments from the DreamerV3 algorithm, including the introduction of a world model regulariser and scaling of returns in the actor critic training step. We find that these developments enable mitigating the exploration challenge with a fixed entropy regulariser and are able to produce strong results without the addition of an exploration-oriented learning system such as Plan2Explore (Section S.LABEL:S-sec:task_agnostic). Avoiding the need to train and run inference on an ensemble of predictors also provides benefits in wall time.
3 Experiments
We perform an empirical study to evaluate the efficacy of WMAR for CL in a set of challenging environments from OpenAI Procgen and Atari Cobbe et al. (2020); Bellemare et al. (2013), which we select for the unique challenges they present in a CL setting, where environments may be extremely similar (Procgen) or dissimilar (Atari). As measures of baseline performance, we compare the performance of WMAR with the most similar model-free and model-based RL algorithms on individual tasks. We choose proximal policy optimisation (PPO) Schulman et al. (2017); Raffin et al. (2021) for the model-free RL algorithm, and use the CNN-LSTM network architecture as it most similarly resembles the architecture of WMAR. We use DreamerV3 as the model-based benchmark.
To measure the efficacy of the long-term distribution matching buffer, we compare CL metrics for WMAR and WMAR with only the FIFO buffer and compare them both with DreamerV3 on Atari. Finally, we compare the performance difference of WMAR when applied to multiple sequential tasks versus when applied to only single tasks. All WMAR experiments are trained on one Nvidia A40 40GB GPU each with all single task benchmarks running within 0.25 days and continual learning benchmarks running within 1 day of wall time, making these experiments widely available and reproducible across research labs.
For more details on experimental parameters and execution time, see Section S.LABEL:S-sec:extime.
3.1 Evaluation
We evaluate the performance by normalising episodic reward to the performance provided by the world model agent’s baseline, and that of a random agent. This enables a normalised comparison in order to measure the relative performance during CL. These results also enable a natural evaluation of forgetting and forward transfer for all environments in each CL suite.
For each CL experiment, we define an ordered suite of tasks $\boldsymbol{\mathcal{T}}=(\tau_{1},\tau_{2},\dots,\tau_{T})$. We define the performance of the agent on task $\tau\in\boldsymbol{\mathcal{T}}$ after $N$ environment steps as $p^{\boldsymbol{\mathcal{T}}}_{\tau}(N)$.
3.1.1 Baseline Normalisation
We perform 2 evaluations of each environment to be considered as ‘baseline’ performance. For each task in the experimental suite $\tau\in\boldsymbol{\mathcal{T}}$, we find the average episodic reward of a random policy $p^{(\tau)}_{\tau}(0)$ and the average episodic reward of the agent on only task $\tau$ after $N$ environment steps $p^{(\tau)}_{\tau}(N)$. The normalized score of subsequent agents will be such that a score of 00 is equal to random performance, and a score of $1$ is equal to the final performance.
$q^{\boldsymbol{\mathcal{T}}}_{\tau}(N)=\frac{p^{\boldsymbol{\mathcal{T}}}_{\tau}(N)-p^{(\tau)}_{\tau}(0)}{p^{(\tau)}_{\tau}(N)-p^{(\tau)}_{\tau}(0)}$ (1)
Average random performance and average final performance of WMAR, when trained on a single task, are used for score normalisation. A score equal to the random performance would correspond to 0, and a score corresponding to the average final performance would correspond to 1.
3.1.2 Continual Learning Assessment
We follow Kessler et al. (2022) in evaluating the efficacy of a continual learning method in forgetting, measuring stability and backward transfer, and forward transfer, also acting as a measure for plasticity. These measurements are recorded for each environment in the experiment suite $\boldsymbol{\mathcal{T}}=(\tau_{1},\tau_{2},\dots,\tau_{T})$. We train the agent on each task for $N$ environment steps. As such, the reference score for task $\tau_{i}$ would be given by $q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(i\times N)$.
Forgetting
The average forgetting for each task is the difference between its performance at the end of training all tasks and its reference score.
$F=\frac{1}{T}\sum_{i=1}^{N}q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(T\times N)-q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}(i\times N)$ (2)
A lower value for forgetting is indicative of improved stability and a superior continual learning method. A negative value for forgetting would imply that the agent has managed to gain performance on earlier tasks, thus exhibiting backward transfer.
Forward transfer
The forward transfer for each task is the expected ratio of average normalised scores between its performance in the continual learning and single task experiments.
$FT$ $=\frac{1}{T}\sum_{i=1}^{T}\frac{S^{\boldsymbol{\mathcal{T}}}_{i}-S^{(\tau_{i})}}{S^{(\tau_{i})}}$ (3)
where
$S^{\boldsymbol{\mathcal{T}}}_{i}$ $=\frac{1}{N}\sum_{j=1}^{N}q_{\tau_{i}}^{\boldsymbol{\mathcal{T}}}((i-1)\times N+j)$ (4)
$S^{(\tau_{i})}$ $=\frac{1}{N}\sum_{j=1}^{N}q_{\tau_{i}}^{(\tau_{i})}(j)$ (5)
The larger the forward transfer, the better the continual learning method. A positive value for forward transfer would imply that the system effectively uses learnt knowledge from previous environments and has accelerated learning in the current environment as a result. When each task is not related to each other, no positive forward transfer may be expected. In this case, a forward transfer of 0 would represent optimal plasticity, and negative values would indicate a barrier to learning newer tasks by previous tasks.
3.2 Environments
We explore the performance of WMAR within the context of transfer between tasks with and without shared structure, which we define as two tasks having similar environment state and observations, action dynamics, and rewards. To investigate the performance on tasks with shared structure, we use OpenAI Procgen’s CoinRun environment and apply visual perturbations to distinguish between different environments. For evaluating the performance on tasks without shared structure, we use 4 different tasks from Atari.
4 Results
4.1 Tasks with Shared Structure
The CoinRun environment provides a number of simple visual perturbations. While the environment’s mechanics remain identical, the environments share both subtle and significant visual differences, which can pose significant risks to CL characteristics. However, we show that WMAR enables constructive interference and positive forward transfer of skills between these environments.
Baseline Performance
We measure the performance of WMAR and other baseline RL algorithms when applied to each task within the suite individually (see Section S.LABEL:S-sec:baseline for parameters and results). The performance of WMAR is approximately 15% better across tasks compared to the PPO (CNN-LSTM) algorithm, but worse compared to DreamerV3, likely due to implementation differences. The ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task.
Continual Learning
The normalised performance is shown in Figure 3 (tabular results in Section LABEL:S-sec:cl). The terminology that denotes the variant of the CoinRun environment is as follows. NB = no background, RT = restricted themes, and MA = Monochrome assets (see Section S.LABEL:S-sec:procgen).
Forgetting and Backward Transfer
Regardless of the choice of replay buffer, the agent is capable of exhibiting desirable forgetting and backward transfer properties. However, we observe a slight benefit to those desiderata when the long-term global distribution matching buffer is used in conjunction with the FIFO buffer, evidenced by the improved forgetting and backward transfer (Table 1). For the FIFO only implementation, no backward transfer on the ‘CoinRun’ environment is observed. This contrasts to the backward transfer observed for the same environment where the distribution matching buffer is present (Section S.LABEL:S-tab:raw_res_coinrun). Further, we observe that the method using the distribution matching buffer shows more consistent performance across tasks with fewer fluctuations in performance during training (Figure 3), which may be a desirable property if there exists a need to snapshot and use the agent at arbitrary points during training.
Forward Transfer
Our results show no evidence of inhibition to the strength of the agent when it is required to learn multiple tasks with shared structure. Rather, we find little difference on the resulting forward transfer regardless of the inclusion of the distribution matching buffer, see Table 1. We find significant forward transfer on all tasks after the first, greatly lowering the amount of time required for a task to reach a certain performance. Importantly, the ordering of the tasks is such that levels in later tasks are either subsets of, or different to levels in earlier tasks. This is so that the agent is unable to completely generalise on earlier tasks just by learning later tasks. Crucially, our results show that strong forward transfer also holds for tasks that are not subsets of earlier tasks.
Figure 3: Continual learning performance of WMAR with (top) or without (bottom) long-term distribution matching buffer, with bold line segments denoting the periods in which certain tasks are being trained. Scores are normalised by Equation 1. The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
Table 1: Average forgetting and forward transfer on tasks with shared structure.
Model Avg. forgetting Avg. fwd. transfer
(Lower is better)
(Higher is better)
WMAR
$\mathbf{-0.07}\pm 0.02$
$\mathbf{0.34}\pm 0.06$
WMAR (FIFO)
$-0.04\pm 0.03$
$0.32\pm 0.06$
4.2 Tasks without Shared Structure
We use a subset of Atari environments in which the world model agent can achieve reasonable performance with little training and follow Machado et al. (2018) in using sticky actions in Atari environments. We run the CL experiment on an arbitrary ordering of the environments as testing each permutation would be prohibitively expensive. We show that significant divergence between environment dynamics in visual differences, reward scale, and game mechanics pose a significant challenge to CL. Moreover, lacking a task oracle, we find predetermined reward scales (Table S.LABEL:S-tab:atari_rew_scales) for each individual environment necessary to enable continual learning.
Baseline Performance
WMAR’s performance remains competitive against other common RL algorithms on single tasks (see Section S.LABEL:S-sec:baseline for parameters and results). Again, the ‘Random’ and ‘WMAR’ scores are used as the normalisation baselines for each task.
Continual Learning
The normalised performance is plotted in Figure 4 (tabular results in Section LABEL:S-sec:cl). It is clear that DreamerV3 is not capable of CL; it only performs well when it is actively training on a given task.
Forgetting
WMAR shows significantly improved stability in learning at all points after the task is learned. With only the FIFO buffer, and in all tasks except the last, the agent loses virtually all performance before the following task has finished training. This is in direct contrast to WMAR with the addition of the long-term global distribution matching buffer where instead, the agent is able to consistently maintain much of its performance on tasks learned earlier on (Table 2). This difference in forgetting emphasises that the distribution matching buffer is necessary to prevent forgetting when tasks are significantly different, as the learning of new tasks causes knowledge of previously learnt tasks to be lost from the world model. As training progresses, samples from the first task have a decreasing probability of being drawn. This is to the point where, by the end of training, samples from the first task have a 12.5% probability of being drawn. Despite this, our results show that WMAR is able to hold much of its performance on the earliest of tasks (Figure 4). Notably, this observation does not apply to all tasks, as we find that the ‘Crazy Climber’ task exhibits the worst overall performance and forgetting even with the distribution matching buffer.
Figure 4: Continual learning performance of WMAR with (top) or without (middle) long-term distribution matching buffer, with bold line segments denoting the periods in which certain tasks are being trained. DreamerV3 benchmark results are also given (bottom). Scores are normalised by Equation 1. The line is the median and the shaded area is between the 25th and 85th percentile, of 5 seeds.
Table 2: Average forgetting and forward transfer on tasks without shared structure.
Model Avg. forgetting Avg. fwd. transfer
(Lower is better)
(Higher is better)
WMAR
$\mathbf{0.10}\pm 0.04$
$-0.35\pm 0.10$
WMAR (FIFO)
$0.68\pm 0.10$
$\mathbf{-0.11}\pm 0.14$
Forward Transfer and Plasticity
We find significant differences in the properties of continual learning on tasks with and without shared structure which are reflected in the results. Contrary to findings in tasks with shared structure, the continual learning setting offers no benefit to forward transfer. Benefits to continual learning in greatly reduced forgetting come at the cost of also reduced forward transfer — effectively trading some plasticity for a significant amount of stability, as the agent is unable to learn later tasks as effectively. We observe a decreasing trend in the agent’s normalized score as it learns more tasks, most apparent in the WMAR experiment, where the agent achieves a normalised score of only $0.49\pm 0.09$ in the ‘Crazy Climber’ environment, substantially lower than the expected 1.0 of the first environment ‘Ms Pacman’ and empirical $0.84\pm 0.04$ of ‘Boxing’ (Table S.LABEL:S-tab:raw_res_atari), which is already well below 1.0. This evidences that it performs worse when learning new tasks after learning unrelated tasks, suggesting that the system is limited by its learning capacity in either the world model or the agent. This limitation thus results in plasticity being gradually lost as training progresses. Moreover, the negative forward transfer shown by experiments with and without the distribution matching buffer suggests that the agent learns newer tasks slower when it has already learned a previous task as compared to if it were randomly initialised, further suggesting that plasticity is a practical constraint in CL with world models.
5 Discussion
WMAR has the ability to perform CL with reduced memory requirements. On tasks with shared structure, it exhibits improved forward and backward transfer, and on tasks without shared structure, it exhibits reduced forgetting. However, fully achieving all desiderata of stability and plasticity under all conditions (with or without shared structure), remains a challenge. Specifically, we show that tasks without shared structure, lacking visual similarity and having different mechanics, may be learned less effectively if they are learned later.
Overall, these results were made possible by taking advantage of a distribution matching buffer running alongside a typical FIFO buffer. This change, coupled with the world model approach to RL, enables improved sample efficiency and can operate with significantly smaller replay buffers. Further, being able to learn tasks quickly benefits scalability of the algorithm.
Our results reiterate findings of world model approaches as strong baselines for CL problems, which we attribute to their natural ability to support off-policy experience replay. Our methods are also largely orthogonal to previous state-of-the-art approaches to combating catastrophic forgetting such as EWC and P&C which work over network parameters, and CLEAR which utilises replay but typically operates over model-free approaches and uses behaviour cloning of the policy from environment input to action output. Yet, we show that WMAR remains capable of eliminating catastrophic forgetting across a diverse set of environments.
5.1 Generalisation of Tasks
While we find tasks with shared structure to enable positive forward and backward transfer, we note that this is not the case in tasks without shared structure. We posit this is due to the reliance on the convolutional feature extractor within the WMAR architecture, enabling it to only take advantage of commonality in the visual domain. Much more abstract similarities or changes in the controls between tasks are likely to negatively impact the forward transfer, causing lower average and maximum performance of later tasks. For example, a change to an environment involving inverting the colours and permuting the controls is likely to induce negative forward transfer. A key distinction between CL with world models and model-free methods is the requirement to reconstruct the high-dimensional observations in the form of images. We posit this induces significant strain on the capacity of the model, possibly in the learning of the latent space, or in the model parameters for encoding and reconstructing the observation, especially if it varies greatly between tasks.
5.2 Reward Scaling without Task Oracle
We also find that approximate reward scaling is required to learn multiple tasks which naturally have extremely different reward scales. While WMAR is sufficiently robust in learning multiple tasks without forgetting even when reward scales are somewhat different, we find this property to not hold when reward scales differ significantly (e.g. by a factor of $10^{2}$), where only the task with the higher reward scale will be learnt. Barring the possibility of access to a task oracle, we believe that this limitation affects all continual learning applications. This limitation is likely a result of the different rewards and subsequent returns causing poorly scaled advantages when training the actor, resulting in the actor only learning tasks with the highest return magnitudes. Experiments involving approaches with automatic scaling of advantages through non-linear squashing transformations proved to hurt learning on individual tasks, so a static, linear, reward transformation was used.
5.3 Limitations of Reservoir Sampling
The WMAR approach brings substantial improvement to memory requirements and CL characteristics, all while significantly reducing catastrophic forgetting. However, the scalability of this approach to a larger number of tasks remains unproven. A key limitation inherited from the use of reservoir sampling is the finite capacity of the long-term global distribution matching buffer. As more tasks are explored and previous tasks are not revisited, an increasing number of samples from previous tasks will inevitably be lost, leading to forgetting. Indeed, we observe some variance in the amount of forgetting to which each task is susceptible. Results from the experiments on tasks without shared structure (Figure 4) suggest latter tasks which may have been learned poorly due to insufficient plasticity in the learning system may experience more forgetting than even earlier tasks.
A plausible approach to improving the performance of WMAR on a larger number of tasks is to increase the size of the long-term global distribution matching buffer. However, this approach naturally inherits limitations from world model approaches and memory-based CL approaches, as doing so incurs significant memory requirements. While techniques such as image compression may alleviate this constraint, RL is frequently done at scale, where the need to send and store a significant volume of image data may pose a significant engineering challenge, reducing the applicability and accessibility of this algorithm.
6 Conclusion and Future Work
We develop a powerful approach for CL within the context of RL approaches using world models. We expect future developments to the concepts shown here, or for existing techniques like behaviour cloning in CLEAR to be combined with our method for further performance gains. Such an adaptation could be countering shifts in the latent distribution as the world model trains and where the actor is frozen. Further, we posit the possibility of tuning the hyperparameters and configuration of WMAR to increase sample efficiency and model capacity to further improve CL. We also propose a difference in learning tasks with and without shared structure, where later tasks may be dissimilar in input, control, and reward scale, and hope for future research to explore the effect of such difference on CL within supervised and non-supervised tasks.